# Addressing the Rank Degeneration in Sequential Recommendation via Singular Spectrum Smoothing

### v1

@@ 3.2. Transformer for SR
where $\mathbf{E}_{v^{-}_{t+1}}^{j}$ denotes the $j$-th negatively sampled item, $N^{-}_{u}$ denotes the number of negatively sampled items from the item set without interaction with the user $u$. The value of $N^{-}_{u}$ affects the recommendation performance significantly (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021). The typical choice in BPR and binary cross-entropy losses use one negatively sampled item $N^{-}_{u}=1$.
@@ 4. Singular Spectrum Smoothing
In this section, we present our proposed SPMRec framework with a novel method based on singular spectrum smoothing, with goals of alleviating the item degeneration and the Transformer sequence rank collapse issues. Specifically, we first discuss the item degeneration and Transformer rank collapse phenomenons. Then we propose a novel spectrum smoothness metric to measure the singular value curve, which indicates how severe are the degeneration and rank collapse. Then we incorporate the smoothness degree into the optimization for both item and sequence sides. We further theoretically connect our dual sides smoothing with the recommendation diversity.
4.1. Item Degeneration and Transformer Rank Collapse
As defined in Eq. (5), the next-item prediction is conducted as the dot product between the sequence output embedding $\mathbf{H}$ and the item embedding matrix $\mathbf{E}$. We argue that the item representations also have the degeneration issue, similar to the word representation degeneration issue discussed in (Gao et al., 2019; Wang et al., 2020; Zhang et al., 2020; Yu et al., 2022). Moreover, we also argue that the sequence output representations suffer from the rank collapse problem because of the degenerated item representations. The item degeneration and Transformer rank collapse show similar behavior, i.e., fast singular value decay, and we argue that item degeneration and Transformer rank collapse are closely related.
4.1.1. Item Degeneration
The item degeneration issue describes the phenomenon that most item embeddings are distributed in a narrow cone, with also the fast singular value decay of the item embedding matrix $\mathbf{E}$. As suggested in (Gao et al., 2019; Yu et al., 2022), the long-tail distribution of words in frequency is the cause of word representations degeneration. In SR, items also share the long-tail distribution. By utilizing the analysis in (Gao et al., 2019), for any tail item $v$, the recommendation loss in Eq. (5) associated with the learnable embedding $\mathbf{E}_{v}$ of the item $v$ is given as:
(6) $\min\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v}\right)+C\right),$
where $C$ includes the constant part without $\mathbf{E}_{v}$. The optimal solution is a vector with the uniformly negative direction of all sequence output embeddings $\mathbf{H}$ and infinite Frobenius norm, i.e., $||\mathbf{E}_{v}||\rightarrow\infty$, using the Theorem 1 in (Gao et al., 2019). With the long-tail distribution of item frequency, most items are cold items. Most cold items degenerate into a narrow space as a cone due to the infinite Frobenius norm. Moreover, as the Frobenius norm of item embedding matrix $||\mathbf{E}||$ is the upper bound of its largest singular value, the infinite Frobenius norm enlarges the largest singular value, causing the fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020), which is also demonstrated in the Fig. (0(b)).
4.1.2. Transformer Rank Collapse
The Transformer rank collapse problem describes the rank degeneration issue of the sequence output embedding $\mathbf{H}$. The Transformer rank degeneration issue shows the fast singular value decay phenomenon, as shown in Fig. (0(a)). Unlike the item degeneration problem, the cause of Transformer rank collapse originates from the architecture design of the Transformer. In (Dong et al., 2021), a theoretical analysis demonstrates that the self-attention module causes the rank collapse exponentially with the network depth if the residual connection is removed. A follow-up work (Zhai et al., 2023) further observes the entropy collapse in the self-attention module. We argue that the rank collapse issue still exists even with the residual connection, which is empirically shown in Fig. (0(a)).
With degenerated item embeddings, we can conclude that the short sequences output embeddings suffer from the rank degeneration. By freezing the item embedding matrix $\mathbf{E}$, we derive the rank collapse of a sequence output embedding $\mathbf{H}_{u}$ for a user $u$ in the recommendation loss as follows:
(7) $\min\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{+}_{t+1}}\right)\right)+\log\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{E}_{v^{-}_{t+1}}^{j}\right),$
where the first component enforces the $\mathbf{H}_{u}$ to be close to the positively interacted items’ representations while the second component pulls the $\mathbf{H}_{u}$ away from negatively sampled items’ representations. When sequences are short, i.e., $|\mathcal{S}^{u}|$ is small, the second component dominates the short sequence representation learning because positively interacted items are much less than negatively sampled items. When the item degeneration happens, i.e., most items have $||\mathbf{E}_{v}||\rightarrow\infty$, the optimal solution of $\mathbf{H}_{u}$ is a vector with negative direction with most items with $||\mathbf{H}||\rightarrow 0$ as more negative items without interactions with the user $u$ are sampled with more epochs of optimization. Using the theories that the Frobenius norm is the upper bound of the largest singular value and the rank is the number of non-zero singular values, $||\mathbf{H}||\rightarrow 0$ indicates the rank losing as other singular values are being suppressed to 0.
@@ 4.2. Degree of Spectrum Smoothness
The observations in Fig. (1) show that both item embeddings and sequence output embeddings have fast singular value decay, which is a crucial indicator of the rank degeneration issue (Wang et al., 2020). It remains unknown how to measure the degree of the rank degeneration issue in the embeddings. We propose a simple, intuitive, and straightforward metric to measure the degree of spectrum smoothness, which is calculated as the Area Under Singular value Curve (AUSC). We first conduct the singular value decomposition (SVD) to decompose both sequence output embeddings $\mathbf{H}$ and item embeddings $\mathbf{E}$ as follows:
(8) $\mathbf{H}=\mathbf{U}_{\mathbf{H}}\Sigma_{\mathbf{H}}\mathbf{V}_{\mathbf{H}}^{\top},\ \ \ \mathbf{E}=\mathbf{U}_{\mathbf{E}}\Sigma_{\mathbf{E}}\mathbf{V}_{\mathbf{E}}^{\top},$
where $\mathbf{U}$ and $\mathbf{V}$ are unitary matrices, and the diagonal matrix $\Sigma$ includes singular values. We represent singular values in the $\Sigma$ as $[\sigma^{\mathbf{H}}_{1},\sigma^{\mathbf{H}}_{2},\dots,\sigma^{\mathbf{H}}_{k}]$ for sequence output embeddings $\mathbf{H}$ and $[\sigma^{\mathbf{E}}_{1},\sigma^{\mathbf{E}}_{2},\dots,\sigma^{\mathbf{E}}_{k}]$ for item embeddings $\mathbf{E}$, where singular values are in descending order as $\sigma_{1}>\sigma_{2}>\dots>\sigma_{k}>0$ and the $\sigma_{1}$ is the largest singular value $\sigma_{\text{max}}$. The area under singular curve (AUSC) is calculated as the sum of normalized singular values as follows:
@@
When the singular value curve decays fast, as observed in the degenerated examples in Fig. (1), the area under the singular curve (AUSC) is typically small because singular values other than the largest one are all small and all singular values are normalized. We can also see that when all singular values approximate the largest singular value, i.e., $|\sigma_{i}-\sigma_{\text{max}}|\approx 0$ for all $\sigma_{i}$, the AUSC is maximized, in which the singular curve becomes flatter. For example, our toy example shown in Fig. (2) compares two singular curves, in which the left one is the fast decaying curve and the right one is an improved and more flat singular curve. The left fast decaying curve has an AUSC value of 13.64, while the right smooth curve has an AUSC value of 26.73.
Figure 2. Intuition of Spectrum Smoothing for addressing item degeneration and sequence rank collapse.
@@ 4.3. Dual Smoothing
We build upon the AUSC to propose a singular spectrum smoothing regularization as a surrogate of optimizing AUSC. With the discussion in Section 4.1, item and sequence rank degeneration issues are intertwined. Thus, we deploy the proposed singular spectrum smoothing regularization to both sequence output and item embeddings, which is dual smoothing. Specifically, the singular spectrum smoothing losses for both sequence $\mathcal{L}_{\text{seq}}$ and item $\mathcal{L}_{\text{item}}$ are conducted as follows:
(10) $\mathcal{L}_{\text{seq}}=-\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}},\ \ \ \ \ \mathcal{L}_{\text{item}}=-\frac{||\mathbf{E}||_{*}}{||\mathbf{E}||_{F}},$
where $\mathbf{H}$ is the sequence output embedding matrix of all users, and $||\mathbf{A}||_{*}$ denotes the nuclear norm as the sum of singular values for any matrix $\mathbf{A}$ ($\mathbf{A}$ can be either $\mathbf{H}$ or $\mathbf{E}$) as follows:
(11) $||\mathbf{A}||_{*}=\text{trace}\left(\left(\mathbf{A}^{\top}\mathbf{A}\right)^{\frac{1}{2}}\right)=\sum_{i}|\sigma_{i}^{\mathbf{A}}|\leq rank(\mathbf{A}),$
and $||\cdot||_{F}$ is the Frobenius norm as the upper bound of the largest singular value, $\sigma_{\text{max}}(\mathbf{A})\leq||\mathbf{A}||_{F}$ (Horn and Johnson, 2012).
Theorem 1 ().
Let $\mathbf{A}$ be a $m\times n$ matrix. Then, the nuclear norm $|A|_{*}$ is a convex relaxation of the rank($\mathbf{A}$). $|\mathbf{A}|_{*}$ provides a lower bound on the rank($\mathbf{A}$).
@@
The singular spectrum smoothing loss is the lower bound of our proposed AUSC smoothness metric as follows:
(12) $\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}}\leq\frac{||\mathbf{H}||_{*}}{\sigma_{\text{max}}^{\mathbf{H}}},\ \ \ \ \frac{||\mathbf{E}||_{*}}{||\mathbf{E}||_{F}}\leq\frac{||\mathbf{E}||_{*}}{\sigma_{\text{max}}^{\mathbf{E}}}.$
@@
The loss minimization of either $\mathcal{L}_{\text{seq}}$ or $\mathcal{L}_{\text{item}}$ achieves the joint effects of maximizing the nuclear norm (encouraging the sum of singular values) and minimizing the Frobenius norm (suppressing the largest singular value) and encourages more non-zero singular values, i.e., preserving the rank. With Theorem 1, the nuclear norm of $\mathbf{H}$ or $\mathbf{E}$ provides a lower bound on the rank (Horn and Johnson, 2012). When $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are minimized, the nuclear norm is maximized, improving the lower bound of the ranks of $\mathbf{H}$ and $\mathbf{E}$. We adopt the Frobenius norm instead of the largest singular value for more stable training. The minimization of $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ maximizes the AUSC of sequence embeddings and item embeddings, i.e., more smooth singular curves are introduced. Specifically, when the singular spectrum smoothing loss is minimized, the Frobenius norm is minimized so that the infinite norm of item embedding $\mathbf{E}$ is alleviated. The largest singular value is suppressed. The computational complexity of singular values calculation can be reduced when $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are conducted in batch samples, which is $O(\text{min}(B^{2}d,Bd^{2}))$ (Cui et al., 2020), and $B$ denotes the batch size and $d$ is the dimension size. Different from (Cui et al., 2021, 2020; Chen et al., 2022), we present the nuclear norm divided by the Frobenius norm on embeddings of sequence and item sides while only nuclear norm of the prediction outputs is maximized in (Cui et al., 2021, 2020; Chen et al., 2022).
4.4. Connection with Recommendation Diversity
In this section, we show the connection between our proposed dual singular spectrum smoothing and recommendation diversity. Specifically, the user-item preference matrix $\mathbf{R}\in\mathbb{R}^{|\mathcal{U}|\times|\mathcal{V}|}$ is given as the dot product between sequence output embedding and item embedding matrix, which is $\mathbf{R}=\mathbf{H}^{\top}\mathbf{E}$.
The rank of $\ \mathbf{R}$ correlates with the recommendation diversity. The rank of the user-item preference prediction matrix $rank(\mathbf{R})$ indicates the number of independent components to represent the user-item preference space (Horn and Johnson, 2012; Chen et al., 2022). The rank of $\ \mathbf{R}$ equals to the rank on the user side and the rank of the item side (Horn and Johnson, 2012). From the user’s perspective, if two users have similar preferences on items, these two users are linearly dependent, and the rank reduces. This happens when two users have similar sequence output embeddings, i.e., the rank collapse of sequence output $\mathbf{H}$. From the item’s perspective, when item embedding matrix $\mathbf{E}$ degenerates, most items are narrowly distributed and hard to distinguish. The rank in the item side can be interpreted as the number of latent item topics to identify items (Koren et al., 2009).
Theorem 2 ().
[…]
With Theorem 2, we can conclude that the rank of $\mathbf{R}$ depends on the rank of sequence embedding $\mathbf{H}$ and the rank of item embeddings $\mathbf{E}$. Our proposed singular spectrum smoothing preserves ranks on both user and item sides. Therefore, the rank of the user-item preference matrix $\mathbf{R}$ is also preserved, and the recommendation diversity is correlated with our proposed smoothing method. We show the empirical relationship between item side spectrum smoothing and the recommendation diversity in Section 5.8.

### v2

@@ 2.3. Next-Item Prediction Optimization
where $\mathbf{M}_{v^{-}_{t+1}}^{j}$ denotes the $j$-th negatively sampled item, $N^{-}_{u}$ denotes the number of negatively sampled items from the item set without interaction with the user $u$. The value of $N^{-}_{u}$ affects the recommendation performance significantly (Ding et al., 2019; Chen et al., 2017; Mao et al., 2021). The typical choice in BPR and binary cross-entropy losses uses one negatively sampled item $N^{-}_{u}=1$.
3. Representation Degeneration and Diversity
3.1. Representation Degeneration Revisit
3.1.1. Revisit of Item Degeneration
Most item embeddings are distributed in a narrow cone when the item degeneration issue occurs, with observation as fast singular value decay of the item embedding matrix $\mathbf{M}$ (Yu et al., 2022; Qiu et al., 2022). As suggested in (Gao et al., 2019; Yu et al., 2022), the long-tail distribution in frequency is the cause of representations degeneration. In SR, items follow the long-tail distribution. By utilizing the analysis in (Gao et al., 2019), for any tail item $v$, the recommendation loss in Eq. (3) associated with the learnable embedding $\mathbf{M}_{v}$ of the item $v$ is given as:
(4) $\min\sum_{\mathcal{S}^{u}\in\mathcal{S}}\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v}\right)+C\right),$
where $C$ includes the constant part without $\mathbf{M}_{v}$. The optimal solution is a vector with the uniformly negative direction of all sequence output embeddings $\mathbf{H}$ and infinite Frobenius norm, i.e., $||\mathbf{M}_{v}||\approx\infty$, using the Theorem 1 in (Gao et al., 2019). With the long-tail distribution of item frequency, most items are cold items. Most cold items degenerate into a narrow space as a cone, with the observation of fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020) and Fig. (2(b)). Moreover, as the Frobenius norm of item embedding matrix $||\mathbf{E}||$ is the upper bound of its largest singular value, the infinite Frobenius norm enlarges the largest singular value, causing the fast singular decay phenomenon observed in (Gao et al., 2019; Wang et al., 2020), which is also demonstrated in the Fig. (2(b)).
3.1.2. Item Degeneration Induces Sequence Degeneration
With degenerated item embeddings, we conclude that output embeddings of short sequences suffer from the degeneration issue. By freezing the item embedding matrix $\mathbf{M}$, we derive the recommendation loss of a sequence output embedding $\mathbf{H}_{u}$ for user $u$ as follows:
(5) $\min\sum_{t=1}^{|\mathcal{S}^{u}|}-\log\left(\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{+}_{t+1}}\right)\right)+\log\sum_{j=1}^{N^{-}_{u}}\exp\left(\mathbf{H}_{(u,t)}^{\top}\mathbf{M}_{v^{-}_{t+1}}^{j}\right),$
where the first component enforces the $\mathbf{H}_{u}$ to be close to the positively interacted items’ representations while the second part pulls the $\mathbf{H}_{u}$ away from negatively sampled items’ representations.
When the item degeneration happens, i.e., most tail items have $||\mathbf{M}_{v}||\approx\infty$, the optimal solution of $\mathbf{H}_{u}$ is a vector in the negative direction with most items, indicating $||\mathbf{H}_{u}||\approx 0$. This is because more negative items are sampled with more optimization steps. $||\mathbf{H}_{u}||\rightarrow 0$ indicates other singular values are being suppressed to 0, based on theories that the Frobenius norm is the upper bound of the largest singular value. When $|\mathcal{S}^{u}|\approx 0$ (i.e., sequences are short), the degeneration issue becomes more severe, as there are much more negatively sampled items than positive items, where the second component dominates the learning process of $\mathbf{H}_{u}$.
3.2. Theoretical Connection with Diversity
3.2.1. Determinant of Embedding as Diversity Measurement
We connect the determinant of embedding with the diversity via the well-defined determinant maximization problem in the determinant point process (DPP) (Kulesza et al., 2012). The determinant maximization problem (Kulesza et al., 2012) has been studied extensively in several machine learning applications, such as core-set discovery (Mahabadi et al., 2019; Mirrokni and Zadimoghaddam, 2015), feature selection (Zadeh et al., 2017), and recommendation re-ranking (Wilhelm et al., 2018; Liu et al., 2022; Chen et al., 2018; Huang et al., 2021). Specifically, the determinant maximization problem is to find a subset from the total item set $\mathcal{Y}\in\mathcal{V}$ with $|\mathcal{Y}|\ll|\mathcal{V}|$. Formally, given the item feature matrix $\mathbf{F}\in\mathbb{R}^{|\mathcal{V}|\times d}$, DPP finds an item subset $\mathcal{Y}$, such that the determinant of the kernel matrix of the item subset $\text{det}(\mathbf{K}_{\mathcal{Y}})$ is maximized as follows:
(6) $\operatorname*{arg\,max}_{\mathcal{Y}\in\mathcal{V}}\text{det}(\mathbf{K}_{\mathcal{Y}}),$
where the kernel $\mathbf{K}_{\mathcal{Y}}(\cdot,\cdot)\in\mathbb{R}^{|\mathcal{Y}|\times|\mathcal{Y}|}$ is typically defined as the dot product of the item features $\mathbf{K}_{\mathcal{Y}}(\mathcal{Y},\mathcal{Y})=\mathbf{F}_{\mathcal{Y}}^{\top}\mathbf{F}_{\mathcal{Y}}$. The existing DPP solutions (Mahabadi et al., 2019; Mirrokni and Zadimoghaddam, 2015; Wilhelm et al., 2018) iteratively expand the diverse subset by searching the item to enlarge the determinant of kernel $\text{det}(\mathbf{K}_{\mathcal{Y}})$.
A subset with two items case (i.e., $\mathcal{Y}=\{v_{i},v_{j}\}$) can be used to demonstrate the relationship between determinant of the kernel $\mathbf{K}_{\mathcal{Y}}$ and the diversity of the subset $\mathcal{Y}$. The determinant of the kernel $\text{det}(\mathbf{K}_{\mathcal{Y}})$ can be extended as $\text{det}(\mathbf{K}_{\mathcal{Y}})=\mathbf{K}_{v_{i}v_{i}}\mathbf{K}_{v_{j}v_{j}}-\mathbf{K}_{v_{i}v_{j}}\mathbf{K}_{v_{j}v_{i}}$, where $\mathbf{K}_{v_{i}v_{j}}\mathbf{K}_{v_{j}v_{i}}$ identifies the correlations with different items. When $\text{det}(\mathbf{K}_{\mathcal{Y}})$ obtains the minimum value 0, $\mathbf{K}_{v_{i}v_{j}}=\mathbf{K}_{v_{i}v_{i}}=\mathbf{K}_{v_{j}v_{j}}$, which indicates that $v_{i}$ and $v_{j}$ are highly correlated, i.e., no diversity. On the contrary, when $\text{det}(\mathbf{K}_{\mathcal{Y}})$ is maximized, the diverse set of items is retrieved.
Extending to the total item set $\mathcal{V}$ for recommendations in SR, the total item set’s diversity can be measured as the determinant of the kernel on $\mathcal{V}$ and the associated item embedding matrix $\mathbf{M}$, which is $\text{det}(\mathbf{K}_{\mathcal{V}})=\text{det}(\mathbf{M}^{\top}\mathbf{M})$. Inductively, given a selected item set with the size as $|\mathcal{Y}|=n$, DPP adds an item $v_{j}$ to the selected item set $\mathcal{Y}$ as $\hat{\mathcal{Y}}$ with size $n+1$. Using Matrix Determinant Lemma, the the determinant of $\hat{\mathcal{Y}}$ is given as:
(7) $\text{det}(\hat{\mathcal{Y}})$ $=\text{det}\left(\begin{bmatrix}\mathbf{K}_{\mathcal{Y}}&\mathbf{K}_{\mathcal{Y},v_{j}}=\mathbf{F}_{\mathcal{Y}}^{\top}\mathbf{F}_{v_{j}}\\ \mathbf{K}_{v_{j},\mathcal{Y}}=\mathbf{F}_{v_{j}}^{\top}\mathbf{F}_{\mathcal{Y}}&\mathbf{K}_{v_{j},v_{j}}=\mathbf{F}_{v_{j}}^{\top}\mathbf{F}_{v_{j}}\end{bmatrix}\right)$
(8) $=\text{det}(\mathbf{K}_{\mathcal{Y}})\cdot\text{det}\left(\mathbf{K}_{v_{j},v_{j}}-\frac{\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}}{\mathbf{K}_{\mathcal{Y}}}\right).$
$\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}$ measures the correlations between the item $v_{j}$ and the selected item set $\mathcal{Y}$. As $\mathcal{Y}$ is selected and fixed, $\text{det}(\mathbf{K}_{\mathcal{Y}})$ and $\mathbf{K}_{\mathcal{Y}}$ can be treated as constants, and $\mathbf{K}_{v_{j},v_{j}}$ is the similarity of the item itself and can also be treated as a constant. DPP aims at finding the $v_{j}$ for maximizing the $\text{det}(\hat{\mathcal{Y}})$, which is equivalent to finding the $v_{j}$ with the least correlations with the selected item set $\mathcal{Y}$:
(9) $\operatorname*{arg\,max}_{v_{j}}\text{det}(\mathbf{K}_{\hat{\mathcal{Y}}})=\operatorname*{arg\,min}_{v_{j}}\mathbf{K}_{v_{j},\mathcal{Y}}\cdot\mathbf{K}_{\mathcal{Y},v_{j}}.$
This inductive argument implies that DPPs favor a diverse item set.
By applying the logorithmic and the SVD decomposition on the embedding matrix $\mathbf{M}$, we obtain the following the relationship between diversity and the singular values of $\mathbf{M}$:
(10) $\log\text{det}(\mathbf{M}^{\top}\mathbf{M})=\log\prod_{i=1}^{d}\left(\sigma^{\mathbf{M}}_{i}\right)^{2}=\sum_{i=1}^{d}2\log\sigma^{\mathbf{M}}_{i}\propto\sum_{i=1}^{d}\sigma^{\mathbf{M}}_{i},$
where $\left[\sigma^{\mathbf{M}}_{1},\sigma^{\mathbf{M}}_{2},\dots,\sigma^{\mathbf{M}}_{d}\right]$ are singular values of the item embedding matrix $\mathbf{M}$, and singular values are in the descending order as $\sigma^{\mathbf{M}}_{1}>\sigma^{\mathbf{M}}_{2}>\dots>\sigma^{\mathbf{M}}_{d}>0$ and the $\sigma^{\mathbf{M}}_{1}$ is the largest singular value $\sigma^{\mathbf{M}}_{\text{max}}$. With Eq. (10), we establish the theoretical relationship between the determinant of embedding and item set diversity. As the 2log only re-scales the singular value, we have shown that the diversity is proportional to the sum of singular values, i.e., the nuclear norm of the embedding $\sum_{i=1}^{d}\sigma^{\mathbf{M}}_{i}=|\mathbf{M}|_{*}$.
Figure 3. Intuition of Singular Spectrum Smoothing for addressing sequence and item representation degeneration.
3.2.2. Representation Degeneration and Insufficient Diversity
The observations in Fig. (2) show that both item embeddings and sequence output embeddings find the fast singular value decay, which is a crucial indicator of the representation degeneration issue (Wang et al., 2020). With SVD to decompose both sequence output embeddings $\mathbf{H}$ and item embeddings $\mathbf{M}$, we obtain singular values of $\mathbf{H}$ and $\mathbf{M}$ as $\left[\sigma^{\mathbf{H}}_{1},\sigma^{\mathbf{H}}_{2},\dots,\sigma^{\mathbf{H}}_{d}\right]$ and $\left[\sigma^{\mathbf{M}}_{1},\sigma^{\mathbf{M}}_{2},\dots,\sigma^{\mathbf{M}}_{d}\right]$.
The fast singular value decay implies that the first singular value is significantly larger than others, i.e., $\sigma^{\mathbf{H}}_{1}\gg\sigma^{\mathbf{H}}_{2}$ and $\sigma^{\mathbf{M}}_{1}\gg\sigma^{\mathbf{M}}_{2}$. It further implies the insufficient diversity of sequence and item embeddings because the determinant (re-scaled sum of singular values) of the embedding, i.e., the diversity from the Eq. (10), achieves a small value when the fast singular value decay happens (most singular values are small values).
@@ 4. Singular Spectrum Smoothing
Motivated by the theoretical connection between the representation degeneration and diversity via the determinant of embedding, we present our proposed SPMRec framework with a novel metric to measure the degree of representation degeneration and a novel singular spectrum smoothing regularization for alleviating the degeneration issue and controlling the diversity simultaneously.
@@ 4.1. Degree of Spectrum Smoothness
We propose a simple, intuitive, and straightforward metric to measure the degree of singular spectrum smoothness, which is calculated as the Area Under Singular value Curve (AUSC). The area under singular curve (AUSC) is calculated as the sum of normalized singular values as follows:
@@
When the singular value curve decays fast, as observed in the degenerated examples in Fig. (2), the area under the singular curve (AUSC) is typically small because singular values other than the largest one are all small and all singular values are normalized. We can also see that when all singular values approximate the largest singular value, i.e., $|\sigma_{i}-\sigma_{\text{max}}|\approx 0$ for all $\sigma_{i}$, the AUSC is maximized, in which the singular curve becomes flatter. For example, our toy example shown in Fig. (3) compares two singular curves, in which the left one is the fast decaying curve and the right one is an improved and more flat singular curve. The left fast decaying curve has an AUSC value of 13.64, while the right smoothed curve has an AUSC value of 26.73.
@@ 4.2. Dual Smoothing
We build upon the AUSC to propose a singular spectrum smoothing regularization as a surrogate of optimizing AUSC. With the discussion in Section 3.1, item and sequence degeneration issues are intertwined. Thus, we deploy the proposed singular spectrum smoothing regularization to both sequence output and item embeddings, which is referred to as dual smoothing. Specifically, the singular spectrum smoothing losses for both sequence $\mathcal{L}_{\text{seq}}$ and item $\mathcal{L}_{\text{item}}$ are conducted as follows:
(12) $\mathcal{L}_{\text{seq}}=-\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}},\ \ \ \ \ \mathcal{L}_{\text{item}}=-\frac{||\mathbf{M}||_{*}}{||\mathbf{M}||_{F}},$
@@
where $\mathbf{H}$ is the sequence output embedding matrix of all users, and The singular spectrum smoothing loss is the lower bound of our proposed AUSC smoothness metric as follows:
(13) $\frac{||\mathbf{H}||_{*}}{||\mathbf{H}||_{F}}\leq\frac{||\mathbf{H}||_{*}}{\sigma_{\text{max}}^{\mathbf{H}}},\ \ \ \ \frac{||\mathbf{M}||_{*}}{||\mathbf{M}||_{F}}\leq\frac{||\mathbf{M}||_{*}}{\sigma_{\text{max}}^{\mathbf{M}}}.$
The loss minimization of either $\mathcal{L}_{\text{seq}}$ or $\mathcal{L}_{\text{item}}$ achieves the joint effects of
  • maximizing the nuclear norm (encouraging the sum of singular values),
  • minimizing the Frobenius norm (suppressing the largest singular value),
  • and enlarges more tail singular values.
@@
When $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are minimized, the nuclear norm is maximized, improving the determinant and diversity of $\mathbf{H}$ and $\mathbf{M}$. We adopt the Frobenius norm instead of the largest singular value for more stable training. The minimization of $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ maximizes the AUSC of sequence embeddings and item embeddings, i.e., more smooth singular curves are introduced. Specifically, when the singular spectrum smoothing loss is minimized, the Frobenius norm is minimized so that the infinite norm of item embedding $\mathbf{M}$ is alleviated. The largest singular value is suppressed.
Computational Efficiency: the computational complexity of singular values calculation can be reduced when $\mathcal{L}_{\text{seq}}$ and $\mathcal{L}_{\text{item}}$ are conducted in batch, which is $O(\text{min}(B^{2}d,Bd^{2}))$ (Cui et al., 2020), and $B$ denotes the batch size and $d$ is the dimension size. Different from (Cui et al., 2021, 2020; Chen et al., 2022), we present the nuclear norm divided by the Frobenius norm on embeddings of sequence and item sides while only nuclear norm of the prediction outputs is maximized in (Cui et al., 2021, 2020; Chen et al., 2022).